Errors in J. Alg. Comb. 1, 133-150. Martin Hildebrand

When referring to irreducible representations λ λ λ, Θ should be used instead of Φ. In equation (2.5), γ should be replaced by δ. On the bottom of p. 145, d_{λ λ λ}^{2} should be replaced with d_{λ̃ λ̃ λ̃}^{2}.

On p. 147 at the end of the line after “By Lemma 5.4, we conclude”, the term α_{n}^{2(n+c)} should be replaced by (α_{n}/q^{i})^{2(n+c)}.

On p. 136, there are two displays where μ μ μ equals something if f = f_{1} and 0 otherwise; in both cases μ μ μ should be replaced by μ μ μ(f).

The author regrets these errors. However, these errors do not affect the results of the paper.